OpenStudy (anonymous):

Find the perimeter of the trapezoid. x/4 + x/4 + x/4 + 2x/4

OpenStudy (lgbasallote):

those are the sides?

OpenStudy (anonymous):

yes

OpenStudy (lgbasallote):

then perimeter equals the sum...where are you having trouble with? you already set up the formula

OpenStudy (anonymous):

I'm just confused on how to add them... :/

OpenStudy (lgbasallote):

hmm okay... \[\frac x4 + \frac x4 + \frac x4 + \frac{2x}{4}\] notice you have the same denominators..if you remember the lesson about that back in your younger years..when you have common denominators then just add the numerators then copy the denominator \[\frac{x + x + x + 2x}{4}\] need any more help?

OpenStudy (anonymous):

oh ok thx for the help :)

OpenStudy (lgbasallote):

you're welcome ^_^